I need to solve the following equation using the graphing method: The perimeter of a soccer field is 340m. The lenght exceeds the width by 50m. I have the the problem solved by using P=2l+2w. How can i solve by using x and Y?
Found 2 solutions by richard1234, ewatrrr:
Answer by richard1234(5390)

If you are talking about 340 = 2x + 2y, the problem won't be any different. Other than the fact that the variables are changed.

Answer by ewatrrr(10682)

 
Hi,
The perimeter of a soccer field is 340m.
340=2l+2w
170 = l + w
Let x and y represent the width and length repectively
170 = y + x OR y = -x + 170 All values > 0 for x and y
satisfying the EQ would work.